OpenStudy (anonymous):

i have a quick question: There is a state soccer tournament with 128 teams competing for 1st place. Each week, half of the teams get eliminated. How many teams remain after 6 weeks? A. 16 B. 8 C. 4 D. 2

OpenStudy (anonymous):

can u plz help me Which of the following sets contains ?

OpenStudy (anonymous):

okayee what can i help you with

OpenStudy (anonymous):

do u kown the Which of the following sets contains ? Answer Rational Numbers Natural Numbers Irrational Numbers Integers

OpenStudy (anonymous):

okayee im nt seeing what sets you're talking bout; The natural (or counting) numbers are 1, 2, 3, 4, 5, etc. There are infinitely many natural numbers. The set of natural numbers is sometimes written N for short. The whole numbers are the natural numbers together with 0. Some textbooks disagree and say the natural numbers include 0. The sum of any two natural numbers is also a natural number (for example, 4 + 2000 = 2004), and the product of any two natural numbers is a natural number (4 × 2000 = 8000). This is not true for subtraction and division, though. The Integers The integers are the set of real numbers consisting of the natural numbers, their additive inverses and zero. The set of integers is sometimes written J or Z for short. The sum, product, and difference of any two integers is also an integer. But this is not true for division... just try 1 ÷ 2. The Rational Numbers The rational numbers are those numbers which can be expressed as a ratio between two integers. For example, the fractions 1/3 and –1111/8 are both rational numbers. All the integers are included in the rational numbers, since any integer z can be written as the ratio z/1. All decimals which terminate are rational numbers (since 8.27 can be written as 827/100.) Decimals which have a repeating pattern after some point are also rationals: for example, 0.083333333... = 1/12. The set of rational numbers is closed under all 4 basic operations, that is, given any two rational numbers, their sum, difference, product, and quotient is also a rational number (as long as we don't divide by 0.) The Irrational Numbers An irrational number is a number that cannot be written as a ratio (or fraction). In decimal form, it never ends or repeats. The ancient Greeks discovered that not all numbers are rational; there are equations that cannot be solved using ratios of integers. The first such equation to be studied was 2 = x2. What number times itself equals 2? is about 1.414, because 1.4142 = 1.999396, which is close to 2. But you'll never hit exactly by squaring a fraction (or terminating decimal). The square root of 2 is an irrational number, meaning its decimal equivalent goes on forever, with no repeating pattern: Historical Note: According to legend, the ancient Greek mathematician who proved that could NOT be written as a ratio of integers p/q made his colleagues so angry that they threw him off a boat and drowned him! Other famous irrational numbers are the golden ratio, a number with great importance to biology: π (pi), the ratio of the cirumference of a circle to its diameter: π = 3.14159265358979... and e, the most important number in calculus: e = 2.71828182845904... Irrational numbers can be further subdivided into algebraic numbers, which are the solutions of some polynomial equation (like and the golden ratio), and transcendental numbers, which are not the solutions of any polynomial equation. π and e are both transcendental. The Real Numbers The real numbers is the set of numbers containing all of the rational numbers and all of the irrational numbers. The real numbers are “all the numbers” on the number line. There are infinitely many real numbers just as there are infinitely many numbers in each of the other sets of numbers. But, it can be proved that the infinity of the real numbers is a bigger infinity. The "smaller", or countable infinity of the integers and rationals is sometimes called 0 (alef-naught), and the uncountable infinity of the reals is called 1 (alef-one). There are even "bigger" infinities, but you should take a set theory class for that! The Complex Numbers The complex numbers are the set {a + bi | a and b are real numbers}, where i is the imaginary unit, –1. (click here for more on imaginary numbers and operations with complex numbers). The complex numbers include the set of real numbers. The real numbers, in the complex system, are written in the form a + 0i = a. a real number. This set is sometimes written as C for short. The set of complex numbers is important because for any polynomial p(x) with real number coefficients, all the solutions of p(x) = 0 will be in C. {FOR FURTHER HELP IM GOING TO NEED TO SEE THE SETS}

OpenStudy (anonymous):

2 teams left after 6 weeks

OpenStudy (anonymous):

people remaining=(1/2)^(number of weeks)*(starting number of teams)

OpenStudy (anonymous):

that's kinda wat i thought but i wasnt sure....Thanks you deserve a medal

OpenStudy (texaschic101):

1st week = 128/2 = 64 2nd week = 64/2 = 32 3rd week = 32/2 = 16 4th week = 16/2 = 8 5th week = 8/2 = 4 6th week = 4/2 = 2 2 teams remain after 6 weeks :)

OpenStudy (anonymous):

thanks @TexasChic101:)))